OpenStudy (deadshot):

What is the simplified form of the quantity y squared plus y minus 6 over the quantity y squared plus 2 y minus 3 ?

OpenStudy (anikhalder):

\[\frac{ y ^{2}+y-6 }{ y ^{2}+2y-3}\]First factorize the numerator separately and then the denominator(middle term factorization)..@DeadShot ..pls tell me if you got it

OpenStudy (deadshot):

The numerator is (x+3)(x-2)

OpenStudy (deadshot):

the denominator is (x+3)(x-1)

OpenStudy (deadshot):

@anikhalder It's \[\frac{ (y+3)(y-2) }{ (y+3)(y-1) }\]so the answer is \[\frac{ (y-2) }{ (y-1) }\] right?

OpenStudy (anikhalder):

Well done 47...:) You are absolutely right

OpenStudy (deadshot):

Alright, Thanks for the help!